OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel?

OpenStudy (anonymous):

I think a not sure

OpenStudy (anonymous):

no i'm not good with math... i can't solve problems like this...

OpenStudy (anonymous):

sorry my mistake you don't need that okay first you need to break it down they gave us the time for each car : car A = 2.4 h car B = 4 h now we need to find the speed so let's call it x car A travelled 22mph faster than B so we can say that : car A = x+22 mph car B = x

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

okay I'm gonna tell you how I would solve it there's a formula distance = speed x time so for car A distance = (X + 22) x 2.4 for car B distance = X x 4 they said that the distance was the same so instead of having an equation for Car A and an equation for Car B seperatly we can join them (X+22) x 2.4 = X x 4

OpenStudy (anonymous):

do you follow so far ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay now solve for X

OpenStudy (anonymous):

x=33?

OpenStudy (anonymous):

yes ! but remember the speed for car A is X + 22

OpenStudy (anonymous):

lol thnx :)